OpenStudy (anonymous):

Could someone explain to me how to do this equation? Solve the system 2x + 2y = -6 and 3x - 2y = 11 by using graph paper or graphing technology. What is the solution to the system? (-1, -7) (1, -4) (2, -1) (3, -2)

OpenStudy (anonymous):

@ganeshie8 @Preetha

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

The first step is: solve both equations for y

OpenStudy (mertsj):

So do that and post your results

OpenStudy (anonymous):

Do i put it in y=mx+b form?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

That would constitute "solving for y"

OpenStudy (anonymous):

y=-x-3 and y=-3x+11/2 ?

OpenStudy (mertsj):

The first one is right. For the second one subtract 3x first and then divide EVERY TERM by -2

OpenStudy (anonymous):

y=-3x/2 +11/2 ?

OpenStudy (mertsj):